OpenStudy (anonymous):

C(x)=2.2+2500/x Evaluate lim x->infinity

zepdrix (zepdrix):

Ummm, think of the graph of 1/x. What happens as you approach larger and larger values of x. It is getting closer and closer to 0. If that is confusing just think about a really large number and compare it to a small number. if x=1, 2500/1 = 2500. As x gets bigger 2500/9999999 is close to 0. So as x gets larger, 2500/x approaches 0. So the limit of your function is 2.2 \c:/

OpenStudy (anonymous):

It says the ans is 2.20. I'm just not sure how to get there. @zepdrix

OpenStudy (anonymous):

The average cost/disc in dollars incurred by Herald Records in pressing x DVDs is given by average cost function C(x)=2.2+2500/x Evaluate lim x->infinity @zepdrix

zepdrix (zepdrix):

Hmm how to explain.. :d the 2500/x becomes less significant as x increases, since the value of 2500/x is getting smaller and smaller. Eventually after enough discs have been made.. the equation is starting to look like this C(x) = 2.2 + 0